A line has a slope of -2/3and a y-intercept of (0, 3). What is the value of y
when x = 4?

Answer by ReadingBoosters(3246)


Use this data to find the linear equation.
point-slope formula is y-y1 = m(x-x1), so
y - 3 = -2/3 (x-0)
y - 3 = -2/3x
y = -2/3x + 3 : this is the linear equation of the line
When x is 4
y = -2/3 (4) + 3
y = -8/3 + 3
y = -8/3 + 3(3/3)
y = -8/3 + 9/3 = ⅓
y = ⅓
(4, ⅓) is the coordinate
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.